Solve the system by the elimination method.
x/2+y/3=7
x/4+2y/3=8

Answer by mananth(16946)

3 x + 2 y = 42 .............1
Total value
3 x + 8 y = 96 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 1
-12 x -8 y = -168
3 x + 8 y = 96
Add the two equations
-9 x = -72
/ -9
x = 8
plug value of x in (1)
3 x + 2 y = 42
24 + 2 y = 42
2 y = 42 -24
2 y = 18
y = 9
x= 8
y= 9
m.ananth@hotmail.ca